A rectangle is a parallelogram with four right angles. A rectangle has a width of 15 feet and a diagonal of a length 22 feet; how long is the rectangle? What is the perimeter of the rectangle? Round to the nearest foot. Show all work to receive full credit.

Length of rectangle:

Perimeter of the rectangle:
: A rectangle is a parallelogram with four right angles. A rectangle has a width of 15 feet and a diagonal of a length 22 feet; how long is the rectangle? What is the perimeter of the rectangle? Round to the nearest foot. Show all work to receive full credit.

Length of rectangle:

Perimeter of the rectangle:

Answer by jojo14344(1031)


Let's see the Rectangle:
drawing(400,400,-12,12,-2,7,line(-10,0,-10,5),line(-10,5,10,5),line(10,5,10,0),line(10,0,-10,0),line(-10,5,10,0),locate(.5,3,22ft=d),locate(-9.9,3,15ft=w))
.
By Pyth Theorem: d^2=l^2+w^2
l^2=d^2-w^2=22^2-15^2=484-225=sqrt(259)
highlight(l=16.09ft)>>>>> 16ft, Length
Then, P=2(l+w)=2(16.09+15)
highlight(P=62.18ft)>>>>>>>62ft
See again the rectangle:
drawing(400,400,-12,12,-2,7,line(-10,0,-10,5),line(-10,5,10,5),line(10,5,10,0),line(10,0,-10,0),line(-10,5,10,0),locate(.5,3,22ft=d),locate(-9.9,3,15ft=w),red(locate(0,.5,16=l)),red(locate(0,5.5,P=62ft)))
Thank you,
Jojo